OpenStudy (1645323):

Find the zero of this equation :x^2 - 3 = -2x

OpenStudy (1645323):

the y= form is x^2+2x-3 and you set it equal to zero but i dont know what to do after

OpenStudy (1645323):

so X^2+2x=3 and i get lost

OpenStudy (anonymous):

\[x^2 + 2x -3 = 0\] now factor them first then solve for x

OpenStudy (1645323):

so is the stuff i posted wrong?

OpenStudy (anonymous):

the first one you post was legit for this problem but you are not done yet you need to factor it

OpenStudy (1645323):

so (x+1)(X+1)?

OpenStudy (anonymous):

no, find two number when multiplied, it gives you -3 and when added, it gives you 2

OpenStudy (1645323):

i dont know

OpenStudy (1645323):

3,-1

OpenStudy (anonymous):

yes, now put that in factored form

OpenStudy (1645323):

(x+3)(X-1)

OpenStudy (1645323):

is that the answer

OpenStudy (anonymous):

not yet the question is find all the zeros so you need to equal each one of the to 0 and solve for x \[(x+3) = 0 \] and \[(x-1) = 0\]

OpenStudy (1645323):

so thats the answer? or is it -3,1?

OpenStudy (anonymous):

correct

OpenStudy (1645323):

which one?

OpenStudy (1645323):

@helpme1.2

OpenStudy (anonymous):

-3,1 are the zeros

OpenStudy (1645323):

ok